OpenStudy (anonymous):

CAN SOMEONE PLEASE HELP ME!!!!!! Simplify. 5 – [–(–2)] Simplify. | –2.8| + | 2.8| Evaluate the expression if d = –3 |d| Evaluate the expression. |18 – 4|

OpenStudy (anonymous):

@CGGURUMANJUNATH Can you help me?

OpenStudy (cggurumanjunath):

1. 3 2. 5.6 3. 3 4. 14

OpenStudy (cggurumanjunath):

5 – [–(–2)] = 5 - (+2) =5-2=3

OpenStudy (cggurumanjunath):

5-2 = 3

OpenStudy (cggurumanjunath):

@iLorita clear ?

OpenStudy (anonymous):

Yes! Thank you! @CGGURUMANJUNATH